I broke my piggy bank again!. This time there was a lot of coins inside. I had nickels, dimes, quarters, twoonies and loonies in the piggy bank. WOW!! I have 4 times as many dimes as I do quarters. I have 3 times as many nickels as I do dimes. I have 5 less loonies than dimes and I have half as many twooines as I do nickels. In total I have 130 coins. How much do I have of each coin and how much money do I have in total?
Answer by stanbon(75887)

I broke my piggy bank again!. This time there was a lot of coins inside. I had nickels, dimes, quarters, twonies and loonies in the piggy bank. WOW!! I have 4 times as many dimes as I do quarters. I have 3 times as many nickels as I do dimes. I have 5 less loonies than dimes and I have half as many twonies as I do nickels. In total I have 130 coins. How much do I have of each coin and how much money do I have in total?
---------------------
Equation::
d = 4q so q = (1/4)d
n = 3d
L = d-5
t = (1/2)n so = (3/2)d
d + q + n + L + t = 130
-------
Substitute and solve for "d"::
d + (d/4) + 3d + d-5 + (3/2)d = 130
-----
5d (7/4)d = 135
20d + 7d = 540
27d = 540
d = 20 (# of dimes)
q = (1/4)d = 5 (# of quarters)
n = 3d = 60 (# of nickels)
L = d-5 = 15 (# of loonies
t = (3/2)d = 30 (# of twonies
-------------
Cheers,
Stan H.
--------------